# Concept Based Explanations and Class Contrasting

### v1

@@ Abstract
Explaining deep neural networks is challenging, due to their large size and non-linearity. In this paper, we introduce a concept-based explanation method, in order to explain the prediction for an individual class, as well as contrasting any two classes, i.e. explain why the model predicts one class over the other. We test it on several openly available classification models trained on ImageNet1K, as well as on a segmentation model trained to detect tumor in stained tissue samples. We perform both qualitative and quantitative tests. For example, for a ResNet50 model from pytorch model zoo, we can use the explanation for why the model predicts a class ’A’ to automatically select six dataset crops where the model does not predict class ’A’. The model then predicts class ’A’ again for the newly combined image in 71% of the cases (works for 710 out of the 1000 classes). [^1] The code including an .ipynb example is available on git: https://github.com/rherdt185/concept-based-explanations-and-class-contrasting
Explainable AI, Computer Vision, Deep Learning, Digital Pathology
@@ 1 Introduction
Interpreting deep neural networks is challenging. At the same time, the range of their use cases increases, also into more critical areas like medicine (Jansen et al., 2023). As the stakes of the application become higher, the need for interpretability increases.
@@
There is a wide range of existing work that focuses on the challenge of explaining such deep neural networks. Earlier methods employ heat maps, where the idea is to highlight pixels in the input that are important for the prediction of the model. Those are usually gradient based (like DeepLift (Shrikumar et al., 2019), Integrated Gradients (Sundararajan et al., 2017), Smooth Grad (Smilkov et al., 2017) or GradCAM (Selvaraju et al., 2017)), or pertubation based (like RISE (Petsiuk et al., 2018)). There are some problems with such heatmap based methods. Previous work raised concerns about the reliability of some gradient based methods (Adebayo et al., 2018) and they may also not increase human understanding of the model (Kim et al., 2018). A more fundamental problem is while those heatmap methods show where the model is seeing something, they do not explain what the model is seeing there, which becomes a problem if what a human sees in the area highlighted by the heatmap does not align with what the model is seeing there (Fel et al., 2023).
@@
As a consequence, other methods emerged that focus on concept based explanation. (Olah et al., 2018) combined saliency maps used in hidden layers with visualizing the internal activations at those layers (therefore combining where the model is seeing something, with what it is seeing there). TCAV (Kim et al., 2018) tests the importance of human defined concepts for the prediction of a specific class (e.g. how important the concept of stripes is for the prediction of the class zebra). ACE (Ghorbani et al., 2019) goes a step further, and extracts those concepts automatically from a dataset, removing the need for human defined concepts. CRAFT (Fel et al., 2023) takes this another step further, and allows to create concept based heatmaps explaining individual images. Further CRAFT uses a human study to show the practical utility of the method.
@@
Figure 1: Example of using our method to explain why a ResNet50 model from pytorch model zoo predicts the class basketball. We extract the six main prototypes for the class basketball (shown in the left column). The middle column shows the same prototypes, but we only sample them from images where the model does not predict basketball (crucially to note, is that for the combined middle column image the model predicts basketball again). The last column shows the original images that the middle column images were cropped out of. They were predicted as ballon, racket, volleyball, volleyball, wooden spoon, rugby ball; But for the model, the combined image from crops of them is predicted as basketball. This works not only for basketball, but for 710 out of the 1000 of the classes.
@@
We also propose a concept based method, both for explaining a single class (like ACE or CRAFT), as well as contrasting any two classes. Unlike ACE as well as CRAFT that both first extract the concepts, and later score them (ACE using TCAV and CRAFT using sobol indice), we first score the activations using attribution and afterwards extract the concepts (so we have no scoring after we extracted the concepts, we assume that all extracted concepts are relevant for the class). And we get a new set of concepts per class. Further, we do not conduct a human user study, but test the faithfulness of the explanation to the model. The idea for the test is that concept based methods extract concepts that are important for the prediction of a given class, and show them to the user in the form of one or more dataset examples (e.g. crops of dataset images).
@@
Then if we extract the n (in the following we always use six) most important concepts for a given class ’A’, then combine the dataset examples for those n concepts and pass the resulting image into the model, it should predict class ’A’ (i.e. we take the concept examples that would be shown to the user as the explanation for class ’A’ and then pass those into the model as a combined image and check whether the model predicts class ’A’ for that combined image). We test this also while taking the dataset examples only from images where the model does not predict class ’A’ (e.g. we take six crops from images where the model does not predict class ’A’, then for a ResNet50 from pytorch model zoo the model predicts class ’A’ for the combined image for 71.0% of the classes). An example of this is shown in Figure 1. If we allow to sample the dataset crops also from images where the model predicts class ’A’, then it works for 94.1% of the classes.
@@ 2 Related Work
This work is focused on providing explanations for already trained deep neural networks, as opposed to training inherently interpretable models (Nauta et al., 2023).
@@
Previous work employed attribution methods, that can be used to generate heat maps, to highlight areas in the input image that are important for the output ((Shrikumar et al., 2019), (Petsiuk et al., 2018)). A fundamental problem with such methods is that they only highlight where in the input the model is seeing something, but do not explain what the model is seeing there. Using such attribution methods at the input layer (like its usually done), can give unreliable results. For some of those methods (Adebayo et al., 2018) has shown that they fail a sanity check and produce similar explanations for a trained and randomly initialized model. Those methods that passed the check, tend to give noisy explanations (Smilkov et al., 2017).
@@
In our case, we use attribution methods (DeepLift with SmoothGrad), but at a hidden layer, not at the input layer. That is similar to (Olah et al., 2018), where attribution methods at a hidden layer were coupled with visualization techniques, therefore answering both where the model is seeing something (attribution) and what it is seeing there (visualization). Similar to (Olah et al., 2018) we also use non-negative matrix factorization (NMF) (Lee & Seung, 1999) on the intermediate activations at a hidden layer to get the explanation down to a manageable number of concepts (we use six). But instead of using all activations, we first filter and score them by attribution, and instead of running NMF on a single image only we run it simultaneously on up to 60 images. Using such a dataset of images to extract relevant concepts is similar to ACE and CRAFT, but we use the whole images to extract concepts instead of using image crops. After we extracted the concepts, then we use image crops to visualize them to the user or run the prediction test on the model.
@@
The test we use for validating the explanation for a single class, is inspired by (Carter et al., 2019). There two images of two classes were combined, to change the prediction of the model to a third class (e.g. combine an image of grey whale with one of a baseball, then the model predicts great white shark for the combined image). Finding such combinations required manual selection of a user based on insights into the model. In our case, finding such combinations works automatically and we use six images instead of two. Also we do not use the whole image, but only a crop of each of the six images (32x32 or 74x74 depending on the hidden layer we use).
Table 1: Post-softmax prediction for the stitched images. Including the desired class in sampling of the patches
Model Layer Average Pred Predicts Desired Class Above 0.1 Above 0.2 Above 0.5
ResNet50 Robust 3.5 0.178 $\pm$ 0.311 22.0% 30.3% 23.4% 16.0%
ResNet50 3.5 0.539 $\pm$ 0.437 56.3% 67.4% 62.2% 54.0%
ResNet34 3.5 0.300 $\pm$ 0.375 34.3% 46.6% 40.6% 28.0%
ResNet50 Robust 4.2 0.415 $\pm$ 0.341 72.8% 74.9% 63.0% 36.4%
ResNet50 4.2 0.898 $\pm$ 0.216 94.1% 98.3% 97.1% 92.1%
ResNet34 4.2 0.852 $\pm$ 0.244 92.5% 97.8% 95.7% 88.4%
Table 2: Post-softmax prediction for the stitched images. Excluding the desired class in sampling of the patches
Model Layer Average Pred Predicts Desired Class Above 0.1 Above 0.2 Above 0.5
ResNet50 Robust 3.5 0.103 $\pm$ 0.229 14.0% 19.4% 15.1% 7.7%
ResNet50 3.5 0.284 $\pm$ 0.389 30.0% 40.0% 35.4% 27.3%
ResNet34 3.5 0.175 $\pm$ 0.300 20.0% 30.1% 24.6% 15.4%
ResNet50 Robust 4.2 0.190 $\pm$ 0.229 42.9% 48.9% 32.2% 11.3%
ResNet50 4.2 0.619 $\pm$ 0.358 71.0% 86.0% 79.0% 62.6%
ResNet34 4.2 0.547 $\pm$ 0.354 65.4% 83.7% 74.3% 55.4%
@@ 3 Models
We evaluate our method on three openly available models trained for classification of natural images (ImageNet1K (Deng et al., 2009)) and an in-house model trained to segment tumor in digital pathology tissue images. For the ImageNet models we use a ResNet50 and a ResNet34 (He et al., 2016) from pytorch model zoo, and a ResNet50 with weights trained for adversarial robustness (Santurkar et al., 2019) (in the following we refer to this model as ResNet50 Robust). For the model trained to segment tumor in digital pathology tissue images we use a modified UNet with a ResNet34 backbone, in the following we refer to this model as ResNet34 Digipath.
In our method we perform operations at a hidden layer of the model. For the digipath model we always use layer3.2 (that would be 21 layers into the ResNet34). For the ImageNet models we use layer3.5 (which would be 27 layers into the ResNet34 or 40 layers into the ResNet50) unless otherwise noted, but we also use layer4.2 (the last convolutional layer) in two experiments. And we always use the layer after relu, so that we have non-negative activations and can use NMF.
@@ 4.1 Generation Process
We want to explain why the model predicts a specific class ’A’, by extracting and visualizing the n (we use six) main features that support that class. Those can be seen as the six prototypes for that class. Figure 2 shows how we extract those six features. First we have a list of images unseen in training (we use the validation data for the imagenet models). We only use images where the model predicts class ’A’ as majority class and we use up to 60 images (if there are less images in the validation data where the model predicts class ’A’, then we use less images). It is important to note here that we ignore the ground truth label, we include only those and all those images where the model predicts class ’A’. Excluding falsely classified images may omit crucial information necessary to understand why the model fails.
For each of the images we extract the intermediate activations at a hidden layer Y (this layer needs to be chosen manually, and that same layer is then used for the whole process). Then we filter and score those activations pixel wise by attribution (we compute the attribution for class ’A’ towards layer Y). As attribution method we always use DeepLift together with SmoothGrad (using 20 samples and a standard deviation in the gaussian noise of 0.25). The attribution is computed from the prediction of the model for class ’A’ before softmax. As baseline for DeepLift we use a white image after normalization for the ImageNet models, and a white image for the digipath model. The attribution method outputs something of shape (batch, channels, height, width), we then compute the mean over the channel dimension to get the contribution per pixel. Then those pixels that have negative attribution are set to zero (we only include those pixels that contribute positively for class ’A’), the others are multiplied by the attribution (so that pixels with higher positive attribution have higher impact).
Then in the third step, those filtered and scored activations are decomposed using non-negative-matrix-factorization (NMF) into n (we use six) features. The NMF outputs a list of embeddings (we do not use those), and n vectors (the basis of the reduced space). In the last step we visualize those six vectors (described in the next section Section 4.2), and interpret those visualizations as showing six prototypes the model is using to predict class ’A’. Those visualizations would be our explanation why the model predicts class ’A’.
@@ 4.2 Visualize NMF Basis Vectors
In this section we describe the method we use to visualize the six NMF basis vectors. Each NMF basis vector, is a vector in a hidden layer of the model. Given such a vector $v$ out of the hidden layer Y, the goal is to visualize $v$ (find an input that represents $v$). We use dataset examples as visualization method.
We sample dataset crops of a size of 32x32 (in case we use layer3.5, if we use layer4.2 then the crop size is 74x74) that have most similarity with $v$. To choose those crops, we first pass all 50000 validation images into the model and access the intermediate activations at layer Y. Then we do a 2x2 average pool on those activations. In the next step we sample the m (we use eight, but that is only for display purposes, for the test we only use the one with the highest cosine similarity) pixels out of those activations that have the highest cosine similarity with $v$. Then we extract the image crops at those positions. For layer3.5, after the 2x2 average pool, each pixel in the activations corresponds to a 32x32 patch in the input image. For layer4.2 that would be 74x74. An example of this is shown in Figure 6. Each row shows the eight most similar dataset crops for one of the NMF basis vectors (the six columns represent the six basis vectors). The similarity decreases from left to right.
Figure 3: Contrast two classes, data generation for the linear classifier, and training the linear classifier
Figure 4: Contrast two classes, explain the difference between two classes in n concepts (n being the number of components in the NMF, we use six) and save NMF-basis
@@ 4.3 Test Explanations
Since we do not know what features the model is using to make its prediction (we do not have ground truths), it is difficult to validate how correct/wrong the explanation is. For our validation test, we first combine the dataset visualization for each of the six concepts into a single image, by concatenating them in the height dimension. Then we check whether the model predicts class ’A’ for that created image. For example, we would pass the first column of Figure 5 (the figure shows the dataset visualizations for Alaskan Malamute) into the model and then check whether/how much the model predicts Alaskan Malamute for that image.
The result for all 1000 classes is shown in Table 1. We run this test for the three ImageNet models. Under average prediction we report the mean and standard deviation of the prediction over running the test for each of the 1000 classes. Predicts desired class reports how often the model predicts the class we wish to explain for the set of prototypes for that class (for the six concepts combined to a single image). Above 0.1 reports how often the model predicts the class we wish to explain with at least 0.1 confidence (score after softmax), similar for above 0.2 and above 0.5. The test works best for the ResNet50 model, when using layer3.5 the model predicts the class we wish to explain in 56.3% of the cases (works for 563 out of the 1000 classes). That still leaves 437 classes where the test fails, for those classes we should be careful about trusting the explanation. When using the last convolutional layer, then it actually works in 94.1% of the cases.
@@
What can be a bit unclear, is whether the model predicts the desired class ’A’ for the visualizations because the visualizations represent the main features the model is using to predict that class ’A’, or whether it is because those visualizations are crops from images where the model is already predicting class ’A’ and then it does the same for the image combined of the crops. Therefore, we also run this test while sampling the visualizations only from images where the model is not predicting class ’A’ as majority class, and those results are shown in Table 2. For example, for the ResNet50 model this test works in 30.0% of the cases (works for 300 out of the 1000 classes) when using layer3.5 and in 71.0% of the cases when using layer4.2 (the last convolutional layer). While this test performs a bit worse compared to allowing class ’A’, it makes a stronger point if it works.
For both tests and all the models, using the last convolutional layer performs better compared to using the earlier layer.

### v2

@@ Abstract
Explaining deep neural networks is challenging, due to their large size and non-linearity. In this paper, we introduce a concept-based explanation method, in order to explain the prediction for an individual class, as well as contrasting any two classes, i.e. explain why the model predicts one class over the other. We test it on several openly available classification models trained on ImageNet1K. We perform both qualitative and quantitative tests. For example, for a ResNet50 model from pytorch model zoo, we can use the explanation for why the model predicts a class ’A’ to automatically select four dataset crops where the model does not predict class ’A’. The model then predicts class ’A’ again for the newly combined image in 91.1% of the cases (works for 911 out of the 1000 classes). [^1] The code including an .ipynb example is available on github:
https://github.com/rherdt185/concept-based-explanations-and-class-contrasting
@@ 1 Introduction
Interpreting deep neural networks is challenging. At the same time, the range of their use cases widens, also into more critical areas like medicine Jansen et al. (2023). As the stakes of the application become higher, interpretability of the behavior of the models becomes more important.
@@
There is a wide range of existing work that focuses on the challenge of explaining such deep neural networks. Earlier methods employ heat maps, where the idea is to highlight pixels in the input that are important for the prediction of the model. Those are usually gradient based (like DeepLift Shrikumar et al. (2019), Integrated Gradients Sundararajan et al. (2017), Smooth Grad Smilkov et al. (2017) or GradCAM Selvaraju et al. (2017)), or pertubation based (like RISE Petsiuk et al. (2018)). There are some problems with such heatmap based methods. Previous work raised concerns about the reliability of some gradient based methods Adebayo et al. (2018) and they may also not increase human understanding of the model Kim et al. (2018). A more fundamental problem is while those heatmap methods show where the model is seeing something, they do not explain what the model is seeing there, which becomes a problem if what a human sees in the area highlighted by the heatmap does not align with what the model is seeing there Fel et al. (2023b).
@@
As a consequence, other methods emerged that focus on concept based explanation. Olah et al. (2018) combined saliency maps used in hidden layers with visualizing the internal activations at those layers (therefore combining where the model is seeing something, with what it is seeing there). TCAV Kim et al. (2018) tests the importance of human defined concepts for the prediction of a specific class (e.g. how important the concept of stripes is for the prediction of the class zebra). ACE Ghorbani et al. (2019) goes a step further, and extracts those concepts automatically from a dataset, removing the need for human defined concepts. CRAFT Fel et al. (2023b) takes this another step further, and allows to create concept based heatmaps explaining individual images. Further CRAFT uses a human study to show the practical utility of the method.
@@
Figure 1: Example of using our method to explain why a ResNet50 model from pytorch model zoo predicts the class basketball. We extract the four main concepts for the class basketball (shown in the left column). The middle column shows the same concepts, but we only sample the crops from images where the model does not predict basketball for the crop and the whole image it came from (crucially to note, is that for the combined middle column image the model predicts basketball again with a confidence of 0.96). The four crops in the middle row were predicted as (from top to bottom): jack-o-lantern, volleyball, volleyball, horizontal bar. But for the model, the combined image from crops of them is predicted as basketball. This works not only for basketball, but for 911 out of the 1000 of the classes. The last column shows the original images that the middle column images were cropped out of. In this instance, even taking the second and third image of the middle row together gets to model to predict basketball with a confidence of 0.87, which points to a bias in the model, since a numbered jersey combined with feet on a hall floor should not be predicted as basketball.
@@
We also propose a concept based method, both for explaining a single class (like ACE or CRAFT), as well as contrasting any two classes. Unlike ACE as well as CRAFT that both first extract the concepts, and later score them (ACE using TCAV and CRAFT using sobol indice), we first score the activations using attribution and afterwards extract the concepts (so we have no scoring after we extracted the concepts, we assume that all extracted concepts are relevant for the class). And we get a new set of concepts per class. Further, we automatically score the explanations using the model itself. The idea for the automatic score is that concept based methods extract concepts that are important for the prediction of a given class, and show them to the user in the form of one or more dataset examples (e.g. crops of dataset images).
@@
Then if we extract the n (4 works good, we tested values between 1 and 10) most important concepts for a given class ’A’, then combine the dataset examples for those n concepts and pass the resulting image into the model, it should predict class ’A’ (i.e. we take the concepts that would be shown to the user as the explanation for class ’A’ and then pass those into the model as a combined image and check whether the model predicts class ’A’ for that combined image). An example of this is shown in Figure 1.
If we explain two similar classes in isolation, then the explanations may also end up being similar and not providing enough information to understand why the model predicts one class over the other. For such a case we propose a method to contrast two classes, providing insight as to why the model predicts one class over the other.
@@ 2 Related Work
This work is focused on providing explanations for already trained deep neural networks, as opposed to training inherently interpretable models Nauta et al. (2023).
@@
Previous work employed attribution methods, that can be used to generate heat maps, to highlight areas in the input image that are important for the output (Shrikumar et al. (2019), Petsiuk et al. (2018)). A fundamental problem with such methods is that they only highlight where in the input the model is seeing something, but do not explain what the model is seeing there. Using such attribution methods at the input layer (like its usually done), can give unreliable results. For some of those methods Adebayo et al. (2018) has shown that they fail a sanity check and produce similar explanations for a trained and randomly initialized model. Those methods that passed the check, tend to give noisy explanations Smilkov et al. (2017).
@@
In our case, we use attribution methods (DeepLift, but we also tested using ’gradient times activation’ Simonyan et al. (2014) and SmoothGrad), but at a hidden layer, not at the input layer. That is similar to Olah et al. (2018), where attribution methods at a hidden layer were coupled with visualization techniques, therefore answering both where the model is seeing something (attribution) and what it is seeing there (visualization). Similar to Olah et al. (2018) we also use non-negative matrix factorization (NMF) Lee and Seung (1999) on the intermediate activations at a hidden layer to get the explanation down to a manageable number of concepts. But instead of using all activations, we first filter and score them by attribution, and instead of running NMF on a single image only we run it simultaneously on 500 images (we tested our method with 50 to 900 images). Using such a dataset of images to extract relevant concepts is similar to ACE and CRAFT, but we use the whole images to extract concepts instead of using image crops. We visualize them using image crops and show them as explanation to the user.
@@
The test we use for validating the explanation for a single class, is inspired by Carter et al. (2019). There two images of two classes were combined, to change the prediction of the model to a third class (e.g. combine an image of grey whale with one of a baseball, then the model predicts great white shark for the combined image). Finding such combinations required manual selection of a user based on insights into the model. In our case, finding such combinations works automatically. Also we do not use the whole image, but stitch crops of images.
Table 1: Prediction test for the explanation. Allowing dataset crops predicted as the target class
Model Layer Average Pred Predicts Target Class
ResNet50 Robust 4.2 0.822 91.0%
ResNet50 4.2 0.985 100.0%
ResNet50 (All 1000 Classes) 4.2 0.981 99.6%
ResNet34 4.2 0.954 97.0%
Table 2: Prediction test for the explanation. Excluding dataset crops predicted as the target class
Model Layer Average Pred Predicts Target Class
ResNet50 Robust 4.2 0.698 86.0%
ResNet50 4.2 $0.884\pm 0.017$ $91.4\%\pm 2.1\%$
ResNet50 (All 1000 Classes) 4.2 0.892 91.1%
ResNet34 4.2 0.850 86.0%
@@ 3 Models
We evaluate our method on three openly available models trained for classification of natural images (ImageNet1K Deng et al. (2009)). We use a ResNet50 and a ResNet34 He et al. (2016) from pytorch model zoo, and a ResNet50 with weights trained for adversarial robustness Santurkar et al. (2019) (in the following we refer to this model as ResNet50 Robust).
In our method we perform operations at a hidden layer of the model. Deeper layers work better, and the penultimate layer (layer4.2) works best, as shown in Figure 4. And we always use the layer after relu, so that we have non-negative activations and can use NMF.
@@ 4.1 Generation Process
We want to explain why the model predicts a specific class ’A’, by extracting and visualizing the n (we usually use 4 for layer4.2, but did a grid search between 1 and 10, see Figure 4) main concepts that support that class. Those can be seen as the n prototypes for that class. Figure 2 shows how we extract those n concepts. First we need a list of images. We only use images where the model predicts class ’A’ as majority class and we usually use 500 images (but using 100 images already performs similar, 50 is a bit worse, see Appendix D for more information about the choice of hyperparameters). It is important to note here that we ignore the ground truth label, we include only those and all those images where the model predicts class ’A’. That is, because excluding falsely classified images may omit crucial information necessary to understand why the model fails.
For each of the images we extract the intermediate activations at a hidden layer Y (this layer needs to be chosen manually, and that same layer is then used for the whole following process). We also investigate which layer works best, coming to the result that deeper layer work better and the penultimate layer works best (similar to the results observed in CRAFT). Then we filter and score those activations pixel wise by attribution (we compute the attribution for class ’A’ towards layer Y). As attribution method we use DeepLift but also tested gradient times activation and SmoothGrad, see Appendix D. The attribution is computed from the prediction of the model for class ’A’ before softmax. As baseline for DeepLift we use a white image after normalization. The attribution method outputs something of shape (batch, channels, height, width), we then compute the mean over the channel dimension to get the contribution per pixel. Then pixels that have negative attribution are set to zero (we only include pixels that contribute positively for class ’A’), the others are multiplied by the attribution (so that pixels with higher positive attribution have higher impact).
Then in the third step, those filtered and scored activations are decomposed using NMF into n concepts. For the NMF decomposition we use scipy Virtanen et al. (2020) version 1.10.1 with the default parameters. The NMF outputs a list of embeddings (we do not use those), and n vectors (the basis of the reduced space). In the last step we visualize those n vectors (described in the next section Section 4.2), and interpret those visualizations as showing n concepts the model is using to predict class ’A’. Those visualizations would be our explanation as to why the model predicts class ’A’.
@@ 4.2 Visualize NMF Basis Vectors
In this section we describe the method we use to visualize the n NMF basis vectors. Each NMF basis vector, is a vector in a hidden layer of the model. Given such a vector $v$ out of the hidden layer Y, the goal is to visualize $v$ (find an input that represents $v$). We use dataset examples as visualization method.
As dataset examples we use use crops of a size of 74x74. Each training image is split into 9 crops in a 3x3 grid, resulting in 11.5 million crops. For each of those crops we compute its spatial average activation in layer Y and then the cosine similarity of the resulting vector with $v$. Then we show those m (usually eight) crops that have the highest cosine similarity.
An example of this is shown in Figure 8. Each row shows the eight most similar dataset crops for one of the NMF basis vectors (the four columns represent the four basis vectors, we used $n=4$ here). The similarity decreases from left to right.
Figure 3: Contrast two classes; Data generation for the linear classifier, and training of the linear classifier
@@ 4.3 Test Explanations
Since we do not know what features the model is using to make its prediction (we do not have ground truths), it is difficult to validate how correct/wrong the explanation is. For our validation test, we first combine the dataset visualization for each of the n concepts into a single image, by concatenating them in the height dimension. Then we check how much the model predicts class ’A’ for that created image and whether it predicts class ’A’ as majority class. For example, we would pass the first column of Figure 6 (the figure shows the dataset visualizations for Alaskan Malamute) into the model and then check whether/how much the model predicts Alaskan Malamute for that image.
The result for 100 classes is shown in Table 1. Due to computational and time constraints we only use every tenth class, resulting in 100 classes (except for the ResNet50 model, there we report the result for every tenth class as well as all 1000 classes). Under average prediction we report the mean prediction over running the test for each of the 100 classes. Predicts target class reports how often the model predicts the target class as majority class. And we use four concepts ($n=4$). The test works best for the ResNet 50 model, When using layer4.2, it works in 99.6% of the cases and it works worst for the ResNet50 Robust model where it works in 91.0% of the cases. In any case where this test fails, the explanation for that class should be discarded (see Section 6 for an example).
@@
What can be a bit unclear, is whether the model predicts the target class ’A’ for the visualizations because the visualizations represent the main features the model is using to predict that class ’A’, or whether it is because the model is already predicting class ’A’ for the individual crops and then it does the same for the combined image. Therefore, we also run this test while only using crops where the model is not predicting class ’A’ as majority class, and those results are shown in Table 2. For example, for the ResNet50 model this test works in 91.1% of the cases when using layer4.2 (the last convolutional layer). While this test performs a bit worse compared to allowing class ’A’, it makes a stronger point if it works.
Figure 4: Testing different number of components in the NMF and different hidden layers. The target class is excluded in sampling of the concepts
In Figure 4 we test what effect of the number of components in the NMF and the chosen hidden layer has. The left plot shows the average prediction for the target class and the right plot shows how often the explanation gets predicted as the target class (would correspond to ’predicts target class’ in Table 2). In both cases higher is better. We also compare to CRAFT. We use CRAFT with the configuration provided in their code example Fel et al. (2023a) (penultimate layer, 10 components in the NMF) and use 500 images as with our method (the code example of CRAFT used 300 images). For our method, we tested four layers (layer1.2, layer2.3, layer3.5 and layer4.2) at different depths of the model. We can see that deeper layer work better, layer1.2 and layer2.3 do not work and the penultimate layer (layer4.2) works best.
When we change the number of components, we need to rerun our method. With CRAFT we only run it once with 10 components, and then use their method to score the importance of each component with relation to the target class to extract the n most important components. Our method slightly outperforms CRAFT when using four or more components and slightly outperforms it overall. The average prediction for CRAFT goes down when using more than 6 components. That might be because the components in CRAFT represent all the activations that are present in the images, and not only those that are predictive for the target class. Therefore, using too many of the components might add features that are not predictive for the class and therefore reduce the prediction.
In the right plot, when using a single component we never match the desired target class (get zero score) by construction, because for visualizing the concepts we exclude those dataset crops where the model predicts the desired target class as majority class. Therefore, when using a single component, we only have a single dataset crop which is never predicted as the target class.
For all other experiments we always use four components (four concepts) for layer4.2 because using more components does not improve the prediction much, and therefore we choose to have fewer concepts.